a painter leans a 20 foot ladder against a wall. The top of the ladder is 16 feet above the ground. How far is the bottom of the ladder from the wall?
Answer by edjones(8007)

a^2+b^2=c^2
a=16, c=20
256+b^2=400
b^2=144
b=12' dist of ladder from wall.
.
Ed